# Scaling Laws for Black-box Adversarial Attacks

### v3

@@ 3.2 Scaling Laws on Image Classifiers
Figure 3: Scaling Laws over model ensemble in a black-box setting. We plot the x-axis on a base-2 logarithmic scale, and we measure both the attack success rate (ASR) and average cross entropy-loss of target models. The cardinality of the model ensembles is varied from $2^{0}$ to $2^{6}$, with models randomly selected for each ensemble. We fit the results of 4 target models by lines for better demonstration.
Before demonstrating scaling laws, we first introduce our experiment settings and hyperparameters.
Surrogate Models. To ensure the transferability and diversity of models, we select 64 models from Torchvision [66] and Timm [84] with different hyperparameters, architectures, and training datasets – Alexnet [51], Densenet [40], Efficientnet [76], Googlenet [74], Inception-V3 [75], Maxvit [80], Mnasnet [77], Mobilenet [39], Regnet [93], Resnet [37], Shufflenet [96], Squeezenet [43], Vgg [72], ViT [27], WideResnet [95], Beit [4], Ghostnet [36], Deit [78], LCnet [14, 85], Repvgg [20], Dpn [11], Mit [90], Bit [49], CvT [88], Twins-SVT [13], and NFNet [6]. The full list of our surrogate models can be found in Appendix B.
Target Models. We carefully choose models that are both widely known and are different from training ensembles in architecture and hyperparameters. We select 7 models as our target models, including Swin-Transformers [61], ConvNeXt [62], and ResNeXt [91].
Dataset. Following previous work [47], we randomly take 100 images from the NIPS17 dataset***https://www.kaggle.com/competitions/nips-2017-non-targeted-adversarial-attack. All images are resized into 224 $\times$ 224. We select 5 categories as target labels, including cat, car, ship, deer and bird.
Hyperparameters. We set the perturbation budget $\epsilon=8/255$ under $\ell_{\infty}$ norm. We take 40 iteration steps of SSA-CWA, while other hyperparameter settings are the same as in [9]. We set the loss function $\mathcal{L}$ to be the cross-entropy loss between the output logit $f(\bm{x})$ and the label $\bm{y}$.
Metrics. We leverage attack success rate (ASR) as our metric, which is (adversarial examples that induce successful attacks /// all adversarial examples).
Experiment Results. The results of ensemble attacks are shown in Fig. 3. From our observations, we empirically derive the scaling laws between the attack success rate (ASR) and ensemble cardinality $T$ as follows:
$\text{ASR}=\alpha\log T+C.$ (9)
It can be explained by the following points:
  • $C$ is influenced by various factors, such as hyperparameters, model architectures, and adversarial targets, and therefore takes different values under different conditions.
  • For a fixed adversarial target, $\alpha$ depends solely on the model itself. Models with the same architecture share the same $\alpha$, regardless of hyperparameter settings.
  • When $T$ is small, the measured ASR exhibits large variance due to different model selections, suggesting that our scaling laws tend to be a statistical result.
  • ASR acts as a good representation of the empirical loss $\mathbb{L}$ defined in Eq. (2). It can be observed from Fig. 3 that there exists strong symmetry between $\mathbb{L}$ and ASR.
  • The scaling laws become less useful as the ASR approaches 1, since ASR no longer reflects $\mathbb{L}$ at that point. This indicates that our observations hold only when the success rate is below the converging threshold.
Our experiment results clearly demonstrate the universality of scaling laws, independent of specific test models or adversarial targets. It is important to note that both our training and evaluation models are open-sourced and pretrained on ImageNet [18], which means they do not strictly satisfy the i.i.d. assumption discussed in Sec. 3.1. Therefore, the theorems derived in Sec. 3.1 should be viewed as a sign of convergence and are not directly related with the empirical result measured by our experiments. Another key improvement observed from our scaling laws is that the evaluation loss $\mathbb{L}$ decreases exponentially with the ensemble cardinality $T$, which is much tighter compared to bounds derived in previous works [41, 94].
3.3 Scaling Laws on Defended Models
Figure 4: Scaling Laws for defended models. While our adversarial examples exhibit high attack success rates across defenses, clear linear scaling laws can be observed from the fitted lines.
Figure 5: Scaling Laws for MLLMs. We measure the ASR of target MLLMs and fit the results with lines. Note that scaling laws still exist, except for the fact that results vary largely between models and targets.
To protect neural networks from attacks, various defense mechanisms have been proposed. Our scaling laws work well with these defense mechanisms, showing that these laws tend to be a general result across defenses.
Target Models. We select six defended models as target models, including 2 adversarially trained models – FGSMAT [52] and Ensemble AT (EnsAT) [79], and 4 input purification defenses – high-level representation guided denoiser (HGD) [56], bit depth reduction (BDR) [24], JPEG compression [16], and DiffPure [67]. We use the same surrogate models and algorithms as Sec. 3.2.
Experiment Results. As demonstrated in Fig. 4, we observe a clear scaling law in all these defended models, further demonstrating its universality. By ensembling a large number of normally trained surrogate models in Sec. 3.2, we improve the attack success rate of DiffPure [67] by more than 30%, adversarially trained models by more than 20%, and high-level guided denoiser [56] by more than 70%. This demonstrates that simply increasing the number of surrogate models can break state-of-the-art defenses like DiffPure [67] and adversarial training [65].
3.4 Scaling Laws on MLLMs
Multimodal Large Language Models (MLLMs) [58, 68, 29] have demonstrated impressive abilities in vision understanding, captioning, and reasoning. In this section, we show that our scaling laws persist in these MLLMs.
Target Models. We still leverage the adversarial examples generated against traditional image classifiers and select 3 popular VLMs – LLaVA-v1.5-7b [58], DeepSeek-VL-7b [17], and Phi-3-vision [29] as target models.
Evaluation Metrics. As MLLMs output text rather than logits, we adopt the definition of ASR from MultiTrust [98], which leverages GPT-4 for determining whether the targeted object is stated in the description. In addition to assessing the correctness of the image classification, it introduces prompts to evaluate the degree of distortion in the crafted image. If GPT determines that the image is highly corrupted, the attack is considered unsuccessful, even if the target model misclassifies the image. This metric ensures that our judgments of successful attacks are of high quality.
Experiment Results. From the results presented in Fig. 5, several key observations can be drawn:
  • All models exhibit a similar value of $C$, indicating that they are all highly robust to attack patterns when the ensemble cardinality $T$ is small.
  • $\alpha$ varies significantly across models, revealing the considerable differences in robustness between MLLMs.
  • Since MLLMs are trained on multimodal data, different adversarial targets significantly influence the attack success rates. This variability may stem from the inherent characteristics of the target labels, as naturally some labels (such as “deer”) are less representative compared to labels like “cat”. This phenomenon, often referred to as class imbalance, has been discussed in prior studies [46].
When compared to image classifiers, MLLMs demonstrate superior robustness across all chosen adversarial targets. This supports the general opinion that multimodal training leads to more robust representations [42], enabling the model to focus more on the essential features of the images rather than specific attack patterns. Furthermore, in contrast to traditional views that scaling model parameters results in increased robustness, we find that among the three target models, Phi-3-vision exhibits the highest robustness. This suggests that the scaling laws derived previously in [5] may not be directly applicable to multimodal models.
@@ 4 Extending Scaling Laws to Vision Encoders
Figure 6: Scaling Laws when we use CLIPs as surrogate models under $\epsilon=8/255$. We take ASR measured by Metric 1 to ensure our results are of high confidence.
In this section, we extend our findings of scaling laws from image classifiers to multimodal foundation models such as CLIP [70]. Sec. 4.1 shows the extension of Eq. (2) to image embeddings, while Sec. 4.2 demonstrates high success rates over commercial multimodal large language models (MLLMs) under our framework.
4.1 Attacking Image Embeddings
Contrastive Language-Image Pretraining (CLIP) is a popular self-supervised framework that can pretrain large-scale language-vision models on web-scale text-image pairs via contrastive learning [70]. Compared to image classifiers that are normally trained on image classification datasets, CLIPs have demonstrated superior zero shot generalization abilities in multimodal downstream tasks [69, 53].
Based on the generalization ability of CLIP models, previous works have utilized them as surrogate models in model ensembles [26, 1, 32]. In fact, the targeted attack framework described in Eq. (2) can be easily extended by replacing the loss function $\mathcal{L}$ with the similarity measure function $\operatorname{sim}(\cdot)$ from CLIP (usually cosine embedding loss in practice), and by substituting the image classifiers $f$ with CLIP’s image encoders $f^{I}$. This modification leads to the following optimization objective:
$\min_{\bm{x}}\frac{1}{T}\sum_{j=1}^{T}\operatorname{sim}(f^{I}_{j}(\bm{x}),f^{I}_{j}(\bm{x}_{\text{tar}})),\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon.$ (10)
Note that besides an empirical loss function, Eq. (10) can also be explained as minimizing the distance between the embeddings of the adversarial examples and the target images on a low-dimension embedding manifold.
4.2 Experiments
@@
We first introduce our settings for attacking pretrained CLIP models and MLLMs used for our evaluation.
@@
Surrogate Models. We select 12 pretrained open-source CLIP models via OpenCLIP [12]. In contrast to previous works that typically use fewer than 4 surrogate models, our ensemble choice is sufficient to yield robust results. The full list of our surrogate clips can be found in Appendix. E.
@@
Target Models. We select 6 commercial MLLMs as our black-box test models - GPT-4o, GPT-4 Turbo [68], Claude-3.5 Sonnet https://www.anthropic.com/news/claude-3-family, Gemini-1.5 Pro https://deepmind.google/technologies/gemini/pro/, Hunyuan [30], Qwen [28] - and 3 open-source MLLMs - LLama-3.2 90B §§§https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/, Pixtral Large https://mistral.ai/news/pixtral-large/, and DeepSeek-VL 7B [17].
ASR Target: Cat Target: Car Target: Ship Target: Deer Target: Bird
$\epsilon$ ($\ell_{\infty}$ norm) $8/255$ $16/255$ $8/255$ $16/255$ $8/255$ $16/255$ $8/255$ $16/255$ $8/255$ $16/255$
GPT-4o 0.58 0.99 0.22 0.52 0.57 0.91 0.40 0.89 0.56 0.97
GPT-4-Turbo 0.68 0.99 0.23 0.55 0.59 0.92 0.38 0.86 0.61 0.97
Claude-3.5 0.56 0.91 0.18 0.42 0.41 0.79 0.37 0.74 0.38 0.80
Gemini-1.5 Pro 0.4 0.92 0.16 0.50 0.35 0.82 0.29 0.82 0.38 0.83
Hunyuan-Vision 0.58 0.95 0.23 0.53 0.59 0.90 0.42 0.88 0.52 0.95
Qwen-VL Plus 0.75 0.98 0.28 0.54 0.65 0.91 0.52 0.92 0.69 0.97
Llama-3.2 90B 0.48 0.88 0.21 0.48 0.42 0.87 0.39 0.87 0.4 0.86
Pixtral Large 0.49 0.88 0.19 0.49 0.46 0.87 0.36 0.77 0.44 0.87
Deepseek-VL 7B 0.54 0.96 0.17 0.52 0.52 0.87 0.44 0.91 0.51 0.93
Table 2: ASR under Metric 2 measured across different settings of $\epsilon$. Remarkably, combined with our methods, most adversarial examples are considered to contain the targets’ features by MLLMs given $\epsilon=16/255$.
Hyperparameters. To demonstrate the scaling laws, we set $\epsilon=8/255$. For practical applications, we follow the common settings in [26, 64] and set $\epsilon=16/255$. All other experiment settings are consistent with those in Sec. 3.
Evaluation Metrics. We employ two types of evaluation metrics in this study. The first metric is the same as in Sec. 3.4. The second metric focuses on the features of adversarial examples, whose goal is to evaluate the insertion of features within adversarial targets into these examples. It requires the target model to generate a description and asks GPT whether the description mentions the adversarial target. Both metrics take ASR as the final score, which we refer to as Metric 1 and Metric 2 in the following sections for simplicity. Test prompts can be found in Appendix D.
Experiment Results. Fig. 6 presents the experiment results on the test models under Metric 1. These results reveal that scaling laws continue to hold when using CLIP models as surrogate models. This outcome seems natural as image features are still correctly learned within the vision encoders. Our method achieves success rates of over 30% on most adversarial targets under Metric 1, indicating that more than $1/3$ of the original images are transformed into high-quality adversarial examples with minimal noise, which demonstrates the efficiency of our method.
We also conduct extensive experiments under Metric 2. From Table. 2 we observe that the high attack success rates of our adversarial examples are generalized results, and the transferability of our adversarial examples is strong enough to fool most popular MLLMs. Compared to previous methods that tried to attack commercial MLLMs by visual adversarial examples [26], our methods exhibit a 40% higher success rate. This demonstrates the strength of our core inspiration of scaling, and highlights the practical applications of our methods. Meanwhile, we find the recently released open-source models Llama-3.2 90B and Pixtral Large exhibiting a better performance than other models generally, which indicates that the AI community is running towards developing more robust, safe and reliable MLLMs.
5 Discussions
In this section, we further demonstrate the scaling laws by using different optimizers, visualizing the adversarial perturbations, and collecting the fitting error.
5.1 Scaling Laws are General Principles
The idea that enlarging the model ensemble induces higher attack success rates has been proposed in previous work [94]. Nevertheless, while prior work usually provides qualitative guesses or bounds, we exploit linear scaling laws by extensive quantitative experiments. While our scaling laws work well with targeted attack settings and the SSA-CWA algorithm, further experiment results reveal that our scaling laws are general principles.

### v4

@@ 3.2 Experimental Setup
Figure 3: Scaling Laws over model ensemble in a black-box setting. We plot the x-axis on a base-2 logarithmic scale, and we measure both the attack success rate (ASR) and average cross entropy-loss of target models. The cardinality of the model ensembles is varied from $2^{0}$ to $2^{6}$, with models randomly selected for each ensemble. We fit the results of selected target models by lines for better demonstration.
To investigate the relationship between the number of surrogate models and attack efficacy, we first define our comprehensive experimental setup. This setup is used to generate a single, fixed set of adversarial examples, which are then evaluated across diverse target models to discover and verify our proposed scaling laws.
3.2.1 Adversarial Example Generation
Surrogate Pool. To ensure the transferability and diversity of our attack, we select a large pool of 64 models from Torchvision [torchvision2016] and Timm [rw2019timm]. This pool includes a wide variety of architectures, hyperparameters, and training datasets, such as AlexNet [krizhevsky2012imagenet], DenseNet [huang2017densely], EfficientNet [tan2019efficientnet], ResNet [he2016deep], ViT [dosovitskiy2020image], DeiT [touvron2021training], and many others. The full list is available in Appendix B.
Ensemble Sampling and Attack Algorithm. We generate adversarial examples by attacking ensembles of surrogate models. For each attack, an ensemble is constructed by randomly sampling $T$ models from our 64-model pool. We vary the ensemble cardinality $T$ across a logarithmic scale from $2^{0}$ to $2^{6}$. We use the SSA-CWA attack [chen2023rethinking] for 40 iteration steps, with a perturbation budget of $\epsilon=8/255$ under the $\ell_{\infty}$ norm. The loss function $\mathcal{L}$ is the standard cross-entropy loss. This process yields 7 distinct sets of adversarial examples, one for each value of $T$.
3.2.2 Target Model Families
To test the universality of any observed scaling phenomena, we evaluate our pre-generated adversarial examples against three distinct families of black-box target models:
  • Family 1: Standard Image Classifiers. We select 7 widely-used models, including Swin-Transformers [liu2021swin], ConvNeXt [liu2022convnet], and ResNeXt [xie2017aggregated].
  • Family 2: Defended Models. We select 6 models representing SOTA defenses, including adversarially trained models (FGSMAT [kurakin2016adversarial], EnsAT [tramer2017ensemble]) and purification defenses (HGD [liao2018defenseadversarialattacksusing], BDR [dong2020benchmarking], JPEG, and DiffPure [nie2022diffusionmodelsadversarialpurification]).
  • Family 3: Multimodal Large Language Models. We select 3 popular MLLMs: LLaVA-1.5-7B [liu2023llava], Qwen-2.5-VL-7B [bai2025qwen25vltechnicalreport], and Llama-3.2-11B-Vision***https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/. These are used to further validate the generality of the laws across different data modalities and architectures.
3.2.3 Dataset and Evaluation Metrics
Dataset and Generation Protocol. Following previous work, we leverage the 1000 images from the ImageNet-1K NIPS 2017 dataset [nips-2017-non-targeted-adversarial-attack, yang2022boostingtransferabilitytargetedadversarial]. To ensure the generality of our findings and create a large-scale testbed, we target 8 distinct categories for each source image [zhang2020understandingadversarialexamplesmutual]. This creates a total of $1000\times 8=8000$ unique source-target pairs for each ensemble configuration. As we generate adversarial examples for 7 distinct ensemble cardinalities $T\in\{2^{0},2^{1},\dots,2^{6}\}$, this results in a large-scale dataset of $7\times 8000=56000$ generated adversarial examples.
To ensure our results are statistically robust and not biased by a single ‘lucky’ or ‘unlucky’ random selection of surrogate models, we employ a batched generation protocol. For any given cardinality $T$, we first partition the 1000 source images into 8 disjoint batches of 125 images. Then, for each batch, we randomly sample a new ensemble of $T$ models from our surrogate pool. This newly sampled ensemble is used to craft the 125 adversarial examples (against all 8 target classes) for its assigned batch. This methodology ensures that the final metrics for each $T$ are averaged over 8 distinct, randomly-drawn ensembles, providing a highly robust and generalizable measurement of the scaling effect.
Metrics. For Family 1 and 2, we measure the Attack Success Rate (ASR), which is defined by the ratio between number of successful attacks (misclassified as target label) and total number of attacks. We also record the cross entropy loss between classifiers’ predictions and target labels for Family 1 to demonstrate the validity of using ASR as the measurements of transferability. For Family 3 (MLLMs), which output text, we adopt a simple evaluation paradigm by string matching the target label text (e.g., ‘sea lion’ for target label 150) and the MLLMs’ descriptions. Since MLLMs produce descriptive text rather than precise classification results, we therefore adopt a relaxed keyword map (provided in Appendix D) to more robustly identify any semantic mentions of the target concept within the outputs.
3.3 Log-Linear Scaling Laws in Image Classifiers
3.3.1 Empirical Observation
We begin our analysis by evaluating the generated adversarial examples on the standard image classifiers. As demonstrated in Fig. 3 (top row), we identify a robust empirical phenomenon: the attack success rate (ASR) exhibits strong log-linear scaling laws with the ensemble cardinality $T$.
As shown by the fitted lines, the ASR increases linearly as we increase $T$ on a logarithmic scale. Concurrently, the average cross-entropy loss against the target label (top row of Fig. 3) exhibits a mirrored log-linear decay. This strong symmetric relationship confirms that ASR is a valid and robust proxy for the true optimization objective, as the decreasing loss directly correlates with the increasing success rate. This suggests that the effectiveness of ensemble attacks scales in a highly predictable manner, where each doubling of $T$ yields a consistent improvement in ASR.
3.3.2 Analysis of the Log-Linear Relationship
We empirically analyze the observable characteristics of these log-linear scaling laws:
  • The Intercept (Baseline Vulnerability): The starting point of the line (at $T=1$) represents the model’s baseline vulnerability to single-model transfer attacks. As seen in Fig. 3, this intercept is influenced by various factors, including the target model’s architecture and the specific adversarial target.
  • The Slope (Scaling Efficiency): The slope of the fitted line represents the ‘scaling efficiency’ of the attack. It quantifies how much ASR is gained for each doubling of $T$. Our results suggest this slope is a key characteristic of the target model’s robustness to scaled attacks.
  • Domain of Validity: Our observed scaling laws hold true within a specific domain. When $T$ is small ($2^{0},2^{1}$), the measured ASR shows large variance depending on the specific models randomly selected. The stable log-linear law emerges as a statistical result as $T$ increases. Meanwhile, as ASR approaches 1.0, the metric becomes saturated, and the linear trend naturally flattens. Our observations hold when ASR is below this saturation threshold.
3.4 Verification of the Scaling Laws’ Universality
(a) Scaling laws on Family 2.
(b) Scaling laws on Family 3.
Figure 4: Verification of the log-linear scaling laws’ universality. (a) The trend persists against SOTA defenses. (b) The trend persists for MLLMs. In both settings, we record the ASR averaged across all 8 target classes to show overall model robustness.
Having established the existence of these log-linear scaling laws, we now test their universality. A critical question is whether this trend is a fragile artifact of standard classifiers or a more fundamental principle of transfer attacks. We test this by evaluating their robustness against active defenses (Family 2) and their generality across entirely different architectures and modalities (Family 3).
3.4.1 Robustness Against Defended Models
We first test the adversarial examples against the defended models (Family 2). The key question is whether SOTA defenses can ‘break’ this observed scaling trend.
As demonstrated in Fig. 4(a), the log-linear scaling laws remain robustly intact. While some defenses successfully lower the overall ASR (seen in the lower ASR values across all $T$), they do not eliminate the fundamental, predictable relationship between $\log T$ and ASR.
This is a critical insight: defenses do not break the underlying scaling behavior. Instead, they primarily increase the ‘cost’ of a successful attack (i.e., requiring a larger $T$ to reach the same ASR). This finding demonstrates that scaling the ensemble is a viable path to breaking even SOTA defenses. By leveraging these scaling laws, we improved the ASR of DiffPure [nie2022diffusionmodelsadversarialpurification] by over 30%, adversarially trained models by over 20%, and HGD [liao2018defenseadversarialattacksusing] by over 50%.
3.4.2 Generalization Across Architectures
Finally, we perform the ultimate test of generality: do the adversarial examples, generated only from standard image classifiers, follow the same scaling laws on MLLMs?
Remarkably, as shown in Fig. 4(b), the log-linear scaling laws persist even in these advanced multimodal models. This cross-modal generalization is non-trivial, as the MLLMs possess vastly different architectures and are trained on different data modalities. This result strongly supports the generality of our scaling laws.
While the log-linear form is general, the specific robustness of each model varies. For instance, LLaVA-1.5-7B and Qwen-2.5-VL-7B have similar parameter counts ($\approx 7$B) and both show high robustness to single-model attacks, yet their averaged ASR curves clearly diverge, indicating different vulnerabilities to scaling. This suggests that the relationship between model scale and adversarial robustness is more complex in MLLMs and challenges the simple ‘model scale = robustness’ narrative [bartoldson2024adversarialrobustnesslimitsscalinglaw].
@@ 4 Extending Scaling Laws to Vision Encoders
Figure 5: Verification of the log-linear scaling laws on SOTA MLLMs. We use an ensemble of CLIP models as surrogates, scaling the cardinality $T$ from 1 to 12. All attacks in this figure are generated using practical budget ($\epsilon=16/255$). The ASR is measured using our single, strict LLM-as-Judge metric (defined in Sec. 4.1). The log-linear scaling laws clearly persist.
In this section, we test the generality and power of our discovered scaling laws. Having established the log-linear scaling laws on standard classifiers, we now extend our investigation to a more complex and practical domain: attacking modern MLLMs by targeting their foundational vision encoders, such as CLIP [radford2021learningtransferablevisualmodels]. We hypothesize that the same scaling principles hold when using the embedding-space attack formulation defined in Sec. 2.1.2.
4.1 Experimental Setup
@@
We first introduce our settings for attacking pretrained CLIP models and evaluating them on MLLM targets.
@@
Surrogate Models. We select 12 pretrained open-source CLIP models via OpenCLIP [cherti2023reproducible]. In contrast to previous works that typically use fewer than 4 surrogate models [zhao2023evaluatingadversarialrobustnesslarge], our scaled ensemble (up to $T=12$) is sufficient to yield robust results. The full model list is in Appendix E.
@@
Target Models. We select 7 state-of-the-art MLLMs as our black-box targets, including 4 commercial models (GPT-4o [openai2024gpt4technicalreport], Claude-3.5-Sonnethttps://www.anthropic.com/news/claude-3-family, Gemini-2.5-Pro [comanici2025gemini], Doubao-Seed-1.6https://seed.bytedance.com/en/blog/introduction-to-techniques-used-in-seed1-6) and 3 state-of-the-art open-source models (Llama-3.2-90B-Vision-Instruct§§§https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/, Qwen3-VL-235B-A22B-Instructhttps://qwen.ai/blog?id=99f0335c4ad9ff6153e517418d48535ab6d8afef&from=research.latest-advancements-list, and DeepSeek-VL2 [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels]).
Dataset and Hyperparameters. We use the same SSA-CWA attack protocol as in Sec. 3.2, but with a simplified dataset for this section. We randomly sample 100 images from the NIPS17 dataset with the same 8 distinct target classes, resulting in 800 unique source-target pairs per ensemble configuration. All experiments in this section are conducted using a unified perturbation budget of $\epsilon=16/255$ under the $\ell_{\infty}$ norm, a common and widely applied setting for practical attacks [dong2023robust, luo2024imageworth1000lies].
Evaluation Metric. We employ a single, unified evaluation metric for all MLLM targets. As MLLMs produce generative outputs instead of classification logits, we adopt the LLM-as-Judge framework from MultiTrust [zhang2024benchmarking]. We leverage GPT-4o to semantically evaluate the target model’s description. An attack is considered successful only if GPT-4 judges that (1) the model’s description mentions the adversarial target concept and (2) the description does not mention factors indicating the image is of low quality, noisy, or modified. The prompt of this strict, high-quality metric is provided in Appendix D.
4.2 Generality and Power of Scaling Laws
Our results, conducted on the latest generation of MLLMs, deliver two key findings.
The log-linear scaling laws persist. Fig. 5 presents the ASR results on our 7 SOTA MLLM targets as a function of ensemble cardinality ($T$). Our first key finding is that the log-linear scaling laws, discovered in Sec. 3.3, robustly generalize to this new domain. This outcome confirms that the principle is not an artifact of standard classifiers but a fundamental property of ensemble attacks. Even when attacking the complex, SOTA vision encoders of models, scaling the number of CLIP surrogates ($T$) yields a predictable, linear increase in ASR on a log scale.
Scaling is a powerful and practical weapon. Having verified the scaling laws (Fig. 5), we now examine its practical impact. Table 1 reports the final ASR from our full scaled ensemble ($T=12$). The results are twofold: they confirm the devastating power of scaling while simultaneously revealing a stark hierarchy in SOTA model robustness. The attack’s power is most evident on Qwen3-VL-235B and GPT-4o, which succumb to a 90.0% and 85.0% average ASR, respectively. This performance represents a significant leap in practical capability, surpassing the 40-50% ASRs reported by previous methods on commercial models [dong2023robust]. However, this effectiveness is not universal. Claude-3.5-Sonnet exhibits exceptional robustness, neutralizing the attack to a mere 31.0% ASR. This resilience stands in sharp contrast to its direct competitor, Gemini-2.5-Pro, which is significantly more vulnerable at 69.0% ASR. The open-source landscape mirrors this variability, with DeepSeek-VL2 (59.0%) and Llama-3.2-90B (58.0%) settling into a moderate robustness tier. These results demonstrate that scaling is a powerful and practical tool, and the resulting ASRs serve as a clear benchmark for the diverging robustness of modern MLLMs.
Target Model - Label Sea Lion Barometer Swing Pickelhaube Fig Lock Computer Snake Average
   GPT-4o 0.86 0.76 0.73 0.92 0.84 0.86 0.93 0.92 0.85
   Claude-3.5-Sonnet 0.25 0.19 0.34 0.38 0.38 0.14 0.38 0.41 0.31
   Gemini-2.5-Pro 0.72 0.44 0.57 0.79 0.86 0.65 0.78 0.68 0.69
   Doubao-Seed-1.6 0.61 0.73 0.76 0.87 0.78 0.70 0.69 0.62 0.72
   Llama-3.2-90B-Vision 0.54 0.45 0.65 0.70 0.54 0.55 0.61 0.61 0.58
   Qwen3-VL-235B 0.90 0.93 0.80 0.95 0.95 0.91 0.98 0.79 0.90
   DeepSeek-VL2 0.68 0.35 0.63 0.50 0.63 0.67 0.54 0.68 0.59
Table 1: Attack Success Rate (ASR) on SOTA MLLMs using our full scaled ensemble ($T=12$) and practical budget ($\epsilon=16/255$). The ASR is measured using our unified LLM-as-Judge metric. We report the per-target ASR and the average ASR across all 8 target classes.
5 Further Analysis and Discussion
In this section, we further demonstrate the robustness of our discovered scaling laws through ablation studies (Sec. 5.1) and provide a deeper analysis into the semantic nature of the perturbations generated via scaling (Sec. 5.2).
5.1 Ablations on the Scaling Laws’ Generality
While our main experiments (Sec. 3) focused on targeted attacks using SSA-CWA, here we verify that the log-linear scaling laws are general principles that hold under different attack settings and across various optimizers.